40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 15 01:17:10 2011, Keiko, Update, LSC, MICH locked and attempt to lock PRCL 
    Reply  Thu Sep 15 02:18:19 2011, Anamaria, Update, LSC, MICH locked and attempt to lock PRCL 
Message ID: 5414     Entry time: Thu Sep 15 02:18:19 2011     In reply to: 5413
Author: Anamaria 
Type: Update 
Category: LSC 
Subject: MICH locked and attempt to lock PRCL 

Kiwamu, Keiko, Anamaria

 

We were able to lock PRC using REFL11I after improving the MICH dark fringe a bit (moving BS) and rotating AS55 and REFL11 such that the signal was maximized in the phases we were using. The dark port is not so dark... but the lock is stable.

I had finished the whole REFL path alignment, but I didn't have a good input beam reference at the time, which is why we had to realign the PDs and the camera. We only had strength to realign 11 and 55. Otherwise, we just need to tweak and center beam on 33 and 165, figure out what's up with 55 and be done with the AP table mods. I hope.

 

Quote:

 Anamaria, Keiko

- We aligned MICH and were successfully locked MICH using AS55Q. The other mirrors were misaligned so that the other degrees of freedom didn't exist. AS55 was fed back to BS. The f2a filters on BS suspension were required to lock, because the pos feedback was unbalanced to angle degrees of freedom.

- We tried to lock PRCL next, however, because we aligned the MICH and the REFL beam paths were changed, REFL PDs didn't have the light anymore. The REFL paths were modified now, so we will try the PRCL locking next.

- We couldn't confirm REFL55 signals although we alined the REFL paths to REFL55 PD.

 

ELOG V3.1.3-